### v1

# Designing ReLU Generative Networks to Enumerate Trees with a Given Tree Edit Distance

Abstract
The generation of trees with a specified tree edit distance has significant applications across various fields, including computational biology, structured data analysis, and image processing. Recently, generative networks have been increasingly employed to synthesize new data that closely resembles the original datasets. However, the appropriate size and depth of generative networks required to generate data with a specified tree edit distance remain unclear. In this paper, we theoretically establish the existence and construction of generative networks capable of producing trees similar to a given tree with respect to the tree edit distance. Specifically, for a given rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$, we prove that all rooted, ordered, and vertex-labeled trees over $\Sigma$ with tree edit distance at most $d$ from $T$ can be generated using a ReLU-based generative network with size $\mathcal{O}(n^{3})$ and constant depth. The proposed networks were implemented and evaluated for generating trees with up to 21 nodes. Due to their deterministic architecture, the networks successfully generated all valid trees within the specified tree edit distance. In contrast, state-of-the-art graph generative models GraphRNN and GraphGDP, which rely on non-deterministic mechanisms, produced significantly fewer valid trees, achieving validation rates of only up to 35% and 48%, respectively. These findings provide a theoretical foundation towards construction of compact generative models and open new directions for exact and valid tree-structured data generation. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
Keywords: Generative networks; ReLU function; trees; tree edit distance; enumeration; Euler string
1 Introduction
Over the past few years, generative networks have been widely studied due to its vast applications in different fields such as natural language processing, data augmentation, DNA sequence synthesis, and drug discovery [1, 2, 3, 4]. Generative networks are a class of machine learning models that learn the underlying patterns, structures, and dependencies in training data. By capturing this statistical information, they can generate new data samples that resemble the original data. These models are not limited to data synthesis but are also used in tasks such as data augmentation, imputation, and representation learning. Applications include generating realistic images, text, audio, and other complex data modalities [5, 6, 7]. For example, the application of generative models in bioinformatics includes motif discovery, secondary structure prediction, drug discovery, cancer research, the generation of new molecules and the analysis of single-cell RNA sequencing data [8], [9], [10], [11].
There are various types of generative models, each with distinct characteristics. Autoencoder-based models include variational autoencoders (VAEs) [12] and denoising autoencoders (DAEs) [13], which are designed to learn compact representations of data by encoding and decoding it. Generative adversarial networks (GANs) [14], including specialized versions like deep convolutional generative adversarial networks (DCGANs) [15], use adversarial training to generate realistic data by having a generator and discriminator compete against each other. Deep belief networks, such as deep Boltzmann machines (DBMs) [16], are probabilistic models that represent complex data distributions through a stack of restricted Boltzmann machines. Generative stochastic networks (GSNs) [17] use stochastic processes to generate data by iteratively refining it. Autoregressive models, including pixel convolutional neural networks (PixelCNN) [18] and pixel recurrent neural networks (PixelRNN) [19], model the distribution of image pixels, generating data pixel by pixel in a sequential manner. The deep recurrent attentive writer (DRAW) model [20] combines recurrent neural networks and attention mechanisms to generate images, focusing on specific parts of the data during the generation process. Diffusion models have recently gained significant popularity in generative AI. These models generate data by first learning how to gradually add noise to real data until it becomes random noise. Then, they are trained to reverse this process, step by step, by removing the noise and recovering the original data distribution. During generation, the model starts with pure noise and progressively denoises it to produce a realistic sample, like an image or audio clip. This step-by-step denoising is guided by a neural network, often trained to predict the noise added at each stage [21], [22].
Selecting the right function family and network model is essential in machine learning. A function family that is too broad may cause high computational costs and overfitting, while a limited one might not produce accurate predictions [23]. Choosing the appropriate network remains challenging, as there is no clear choice for every problem. The universal approximation theorem states that a two-layer neural network can approximate any Borel measurable function [24]. But such networks often require a large number of nodes. Studies have examined how the choice of function family relates to network size, revealing that deeper networks significantly enhance representational power [25, 26]. Not all functions can be efficiently represented by any architecture. For instance, Telgarsky [27] identified functions requiring exponentially more nodes in shallow networks compared to deep ones. Szymanski and McCane [28] showed that deep networks are well-suited for modeling periodic functions, while Chatziafratis et al. [29] established width lower bounds based on depth for such functions. Hanin and Rolnick [30] further found that networks with piecewise linear activation functions do not exponentially increase their expressive regions. Additionally, Bengio et al.[31] and Biau et al.[32] demonstrated that decision trees and random forests can be approximated by neural networks using sigmoidal, Heaviside, or tanh activation functions. Kumano and Akutsu [23] later extended this result to networks using ReLU and similar activations. Recently, Ghafoor and Akutsu [33] discussed the existence of generative networks with ReLU as activation function and constant depth to generate similar strings with a given edit distance.
Selkow [34] introduced the problem of tree edit distance as a generalization of the classical string edit distance problem. The tree edit distance problem has several applications in applied fields including computational biology [35, 36, 37, 38], analysis of structured data [39, 40, 41], and image processing [42, 43, 44, 45]. Different algorithms have been developed to compute the tree edit distance between two rooted, labeled and ordered trees. For instance, Tai [46] proposed an algorithm for the tree edit problem with time complexity $\mathcal{O}(n^{6})$, where $n$ is the size of the underlying tree. Zhang and Shasha [47], Klein [48] and Demaine et al. [49] proposed improved algorithms with a time complexity $\mathcal{O}(n^{4})$, $\mathcal{O}({n^{3}}\log n)$ and $O(n^{3})$, respectively. Later on Bringmann et al. [50] further improved the complexity to $\mathcal{O}(n^{3-\epsilon})$ for weighted trees. In 2022, Mao [51] introduced an algorithm for unweighted trees with complexity $\mathcal{O}(n^{2.9148})$. Recently, Nogler et al. [52], proposed an efficient algorithm with complexity $\mathcal{O}(n^{3/2\Omega(\sqrt{\log n})})$ for weighted trees and $\mathcal{O}(n^{2.6857})$ for unweighted trees.
Recent years have witnessed significant advancements in structured generative networks, with a growing focus on models that balance empirical performance and structural validity. GraphRNN is a foundational autoregressive model that generates graphs by sequentially adding nodes and edges, widely used in molecular design [53]. Building on this, Wang et al. [54] proposed a variational autoregressive model that learns generation order dynamically, achieving state-of-the-art molecular graph results without diffusion. AutoGraph [55] applies transformers to autoregressively generate graphs as sequences. TreeGAN by Liu et al. [56] is a syntax-aware generative adversarial network designed for sequence generation that respects tree-structured syntactic constraints. In parallel, diffusion-based generative models have recently advanced structured data generation. For instance, Huang et al. [57] proposed a continuous-time generative diffusion model, GraphGDP, to generate permutation-invariant graphs. Liu et al. [58] introduced a beta-noise process that effectively models both discrete graph structures and continuous node attributes, achieving strong results on biochemical and social network benchmarks. The framework by Madeira et al. [59] enforced hard structural constraints such as planarity or acyclicity via an edge-absorbing noise mechanism, ensuring generated graphs rigorously maintain desired properties throughout the diffusion process. While existing generative models offer powerful empirical frameworks for producing high-quality and diverse structured data, their guarantees are inherently probabilistic and data-dependent. These methods require training on limited datasets, and their performance is influenced by the quality and coverage of this data. As a result, they cannot provide exact enumeration of the underlying combinatorial space, nor can they ensure complete validity or coverage of all possible structured instances.
As a step towards addressing these issues we study the exact and deterministic generation of all rooted, ordered, and vertex-labeled trees similar to a given tree using neural networks and prove the existence of ReLU-based generative networks for tree edit distance. Given a rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from a symbol set $\Sigma$, we theoretically establish the existence of ReLU-activated generative networks capable of producing all rooted trees over $\Sigma$ within tree edit distance at most $d$ from $T$. The key idea of our approach is to first construct a directed, rooted, ordered, and edge-labeled tree based on $T$. We then reduce the tree edit distance problem to a string edit distance problem by representing the tree as an Euler string [48], obtained through a depth-first search (DFS) traversal. The proposed networks are applied on trees with up to 21 nodes to generate similar trees, and are compared with the state-of-the-art graph generative models GraphRNN by You et al. [53] and GraphGDP by Huang et al. [57]. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
The paper is organized as follows: Preliminaries are discussed in Section 2. ReLU generative networks that can identify the indices and labels of the directed edges in the Euler string are discussed in Section 3. Existence of ReLU networks to generate all trees with tree edit distance at most $d$ due to substitution operations is discussed in Section 4. Existence of ReLU networks to generate all trees with edit distance at most (resp., exactly) $d$ due to deletion (resp., insertion) operations is discussed in Section 5. Generation of all trees with tree edit distance at most $d$ due to simultaneous application of deletion, substitution and insertion operations by using a ReLU network is discussed in Section 6. Computational experiments are discussed in Section 7. A conclusion and future directions are given in Section 8. Proofs of some theorems, examples and explanations of the program codes with sample instances are given in Appendix 9.
2 Preliminaries
Edit distance between two vertex-labeled, rooted and ordered trees $T$ and $U$ is defined as the minimum number of operations needed to transform $T$ into $U$. These operations are substitution, deletion, and insertion. Substitution involves simply changing the label of a node in $T$; deletion removes a non-root node $a$ in $T$, reassigning its parent $b$ as the new parent of all children of $a$; and insertion is the complement of the deletion operation, i.e., a node $a$ is inserted as a child of a node $b$, and $a$ is set as the new parent of an ordered subset of consecutive children of $b$. The order among the children is preserved during both the deletion and insertion operations, as illustrated in Fig. 1.
Figure 1: Tree deletion and insertion operations. In $T$, the node with label 7 is the parent of the node with label 5, which has been deleted. The node with label 7 becomes the new parent of the children with labels 1 and 6 of node 5 in $U$. Similarly, a node with label 5 is inserted as a child of the node with label 7 in $U$, and the nodes with labels 1 and 6 are set as children of node 5 in $T$. The order among the children 1 and 6 is preserved in the deletion and insertion operations.
Let $T$ be a vertex-labeled, rooted and ordered tree with $n$ edges ($n+1$ vertices) with vertex labels from the set $\Sigma=\{1,2,\ldots,m\}$, and left-to-right ordering on the siblings of each vertex. We consider the depth-first search (DFS) index on the vertices of $T$ starting from the root with index 0. For $T$, we define a directed edge-labeled, rooted and ordered tree with $n+1$ vertices and $2n$ edges as follows: replace the edge between any two adjacent vertices $u$ and $v$ with labels $a$ and $b$, resp., where $u$ is the parent of $v$, by two directed edges $(u,v)$ and $(v,u)$ with labels $b$ and $b+m$, respectively. In this setting, we call $(u,v)$ and $(v,u)$, the inward edge and the outward edge, resp., of the vertex $v$. If $i$ is the DFS index of $v$, then we call the inward and outward edges of $v$, the inward and outward edges of $i$. We call $(v,u)$, the outward edge of the inward edge $(u,v)$. The Euler string of $T$ is defined to be the string obtained by listing the labels of the inward and outward edges of the directed tree corresponding to $T$ in the DFS order on edges starting from index $1$. We denote by $E(T)=t_{1},t_{2},\ldots,t_{2n}$, the Euler string of $T$ with $n$ edges. An example tree $T$, its directed tree, and Euler string are given in Figs. 2(a), (b), and (c), respectively. Henceforth, we will use the terms, edge and label interchangeably.
Figure 2: (a) A vertex-labeled, rooted and ordered tree $T$ with six vertices, root $r$, label set $\Sigma=\{1,2,3,4,5\}$ and a random sequence $x_{1},x_{2},x_{3}=1,3,0$, where the labels are depicted inside the vertices, and the DFS indices are shown in green; (b) The directed tree corresponding to $T$ given in (a). The inward edges and outward edges are depicted by solid and dashed directed lines, respectively. The labels and DFS indices of these edges are shown in black and brown color, respectively. The vertex $u$ with label $3$ is the parent of $v$ with label 22. Corresponding to the edge $uv$ in $T$, there is an inward edge $(u,v)$ and an outward edge $(v,u)$ with labels $2$ and $7$, respectively, in the directed tree. These edges $(u,v)$ and $(v,u)$ are the inward and outward, resp., edges of $x_{2}=3$ since the DFS index of $v$ is 3 in $T$. Note that there is no edge that corresponds to $x_{3}=0$; and (c) The Euler string $E(T)$.
Observe that a tree can be completely determined by its Euler string, i.e., $E(T)$ is a canonical representation of $T$ when the labels of roots of the underlying trees is fixed. Therefore any tree edit operations (substitution, deletion, and insertion) on a non-root vertex $u$ of a given tree can be viewed as edit operations on the Euler string of the tree on the entries that correspond to the inward edge $(u,v)$ and the outward edge $(v,u)$ with some refinements to obtain the desired tree. Furthermore, a vertex in $T$ can be specified by its DFS index. Therefore in the rest of the paper, we will use a random sequence $x_{1},x_{2},\ldots,x_{d}$, $d\geq 1$, with integers $x_{j}\in[0,n]$, unless stated otherwise, to specify the DFS indices of the vertices under consideration in a tree with $n+1$ vertices. We ignore $x_{j}=0$ and repeated entries in the case of substitution and deletion operations. An example random sequence is given in Fig. 2(a). In the rest of the discussion, we call a vertex-labeled or edge-labeled, rooted and ordered tree simply a tree. A list of symbols, variables, and their descriptions used throughout the discussion is provided in Table 3.
Table 1: List of symbols, variables, and their descriptions.
Symbols Explanations
$T$ Vertex-labeled, rooted and ordered tree. See Fig. 1.
$(u,v)$ Inward edge of $v$, where $u$ is the parent of $v$. See Fig. 2(b.)
$(v,u)$ Outward edge of $v$, where $u$ is the parent of $v$. See Fig. 2(b).
$E(T)$ Euler string of tree $T$. See Fig. 2(c).
$n$ Number of edges in a tree.
$d$ Edit distance.
$B$ A sufficiently large number.
$\Sigma$ Set $\{1,2,\ldots,m\}$ of labels.
$\rm{in}_{\ell i}$ Number of inward edges between $\ell$}-th and $i$-th entries of an Euler string. See Fig. 4 .
$\rm{out}_{\ell i}$ Number of outward edges between $\ell$}-th and $i$-th entries of an Euler string. See Fig. 4.
DFS index Depth first search indexing of vertices. See Fig. 2(a)
TSd Generative ReLU network for tree edit distance $d$ due to substitution. See Fig. 6.
TDd Generative ReLU network for tree edit distance $d$ due to deletion. See Fig. 7.
TId Generative ReLU network for tree edit distance $d$ due to insertion. See Fig. 8.
TEd Generative ReLU network for tree edit distance $d$ due to substitution, deletion or insertion. See Fig. 9.
$x_{1},x_{2},\ldots,x_{2d}$ A random input sequence for TSd. $x_{j},1\leq j\leq d$ specifies the DFS index of a vertex for substitution, and $x_{d+j},1\leq j\leq d$ denote the value to be substituted. See Fig. 6.
$x_{1},x_{2},\ldots,x_{d}$ A random input sequence for TDd. $x_{j}$ specifies the DFS index of a vertex of a tree to be deleted. See Fig. 2(a).
$x_{1},x_{2},\ldots,x_{4d}$ A random input sequence for TId. $x_{j},1\leq j\leq d$ specifies the DFS index of a vertex for insertion, and $x_{d+j}$ and $x_{2d+j},1\leq j\leq d$ specify bounds on the children of the vertex with index $x_{j}$, and $x_{3d+j},1\leq j\leq d$ denotes the value to be inserted. See Fig. 8.
$x_{1},x_{2},\ldots,x_{7d}$ A random input sequence for TEd. $x_{j},1\leq j\leq d$ specifies the input for deletion, $x_{d+j},1\leq j\leq 2d$ specifies the input for substitution, and $x_{3d+j},1\leq j\leq 4d$ specifies the input for insertion. See Fig. 9.
All local variables used in Lemma 1 are explained in Example 1 and Fig. 3.
All local variables used in Proposition 1 are explained in Example 2 and Fig. 4.
All local variables used in Lemma 2 are explained in Example 3 and Fig. 5.
All local variables used in Theorem 1 are explained in Example 4 and Fig. 12.
All local variables used in Theorem 2 are explained in Example 5 and Fig. 13.
All local variables used in Theorem 3 are explained in Example 6 and Fig. 14.
All local variables used in Theorem 4 are explained in Example 7.
3 Identification of Edge Labels by ReLU Network
We focus on designing generative networks with ReLU as an activation function to generate all trees that are similar to a given tree. More precisely, we are interested in the following problem:
Input: A rooted, ordered, and vertex-labeled tree $T$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$.
Output: Construct generative networks with ReLU as an activation function that can generate all rooted, ordered, and vertex-labeled trees over $\Sigma$ with tree edit distance at most $d$ from $T$.
We target this problem by reducing the tree edit distance problem to the string edit distance problem by representing trees as their Euler strings as explained in Section 2. As a sub-task, the positions and labels of the under consideration inward and outward edges in the Euler string are required to perform the edit operations. However such positions and labels are not readily available. Therefore we first discuss the existence of ReLU networks to identify the positions and labels of the inward and outward edges in an Euler string corresponding to a given random sequence $x_{1},x_{2},\ldots,x_{d}$ in Lemma 1.
Lemma 1.
Let $T$ be a tree with $n$ edges, and $x=x_{1},x_{2},\ldots,x_{d}$ be a random DFS sequence of integers over the interval $[0,n]$. Then there exists a ReLU network with size $\mathcal{O}(dn)$ and constant depth that can identify the label of inward edge of the vertex with non-zero DFS index $x_{j}$ in the Euler string $E(T)$.
Proof.
Let $E(T)=t_{1},t_{2},\ldots,t_{2n}$. The following system of equations can be used to obtain the labels of the required inward edges in $E(T)$ (see Example 1), where $i\in\{1,2,\ldots,2n\}$, $j\in\{1,2,\ldots,d\}$ and $C$ is a large number.
$p_{i}$ $=\begin{cases}1&\text{~if~}t_{i}\leq m,\\ 0&\text{~otherwise},\end{cases}$ (1)
$p^{\prime}_{i}$ $=\max(\sum_{k=1}^{i}p_{k}-C\delta(p_{i},0),0),$ (2)
$p^{\prime\prime}_{i}$ $=p^{\prime}_{i}+\max(2n-C(1-\delta(p^{\prime}_{i},0)),0),$ (3)
$q_{ji}$ $=\delta(p^{\prime\prime}_{i},x_{j}),$ (4)
$r^{\prime}_{ji}$ $=t_{i}\cdot q_{ji}.$ (5)
Eq. (1) outputs a binary variable $p_{i}$ which is 1 if and only if the $i$-th entry of $E(T)$ is the label of an inward edge. Eq. (2) identifies the DFS index of only inward edges in $E(T)$ (see Example 1). That is $p_{i}^{\prime}=\ell\neq 0$ if and only if the $i$-th entry of $E(T)$ corresponds to the $\ell$-th inward edge in the directed tree. Eq. (3) replaces $p_{i}^{\prime}=0$ by $2n$ to ignore the root case. The variable $q_{ji}=1$ if and only if $p_{i}^{\prime}=x_{j}$ in Eq. (4), and Eq. (5) is used to identify the labels of the desired inward edges. Note that all these equations involve the maximum function or $\delta$ function which can be simulated by the ReLU activation function based on Proposition 1 by Ghafoor and Akutsu [33]. Therefore we can construct an eight-layer neural network with ReLU as an activation function with size $\mathcal{O}(dn)$ and constant depth that can identify the labels of the inward edges of non-zero $x_{j}$. ∎
A demonstration of Lemma 1 is given in Example 1.
Example 1.
Consider the tree $T$ shown in Fig. (2)(a) with $E(T)=3,2,7,2,$ $4,9,7,4,9,8$, and the random sequence $x=1,3,0$. We wish to identify the labels of the inward edges of $x_{j}\neq 0$ in $E(T)$. For $x_{1}=1$ and $x_{2}=3$, the labels of the inward edges are $3$ and $2$, resp., whereas $x_{3}=0$ does not correspond to any inward edge, and therefore it is ignored. The variables that are used in the process of obtaining the required labels by using Lemma 1 are discussed and illustrated in Table 2 and Fig. (3).
Table 2: The variables, their meaning and example values used in Lemma 1.
Variable          Meaning           Value
$x_{j}$ Specify the inward edge of $x_{j}$ of which the label is required. $x=1,3,0$ (Fig. 2(a))
$p_{i}$ A binary variable which is one if there is an inward edge at the $i$-th position of $E(T)$. $p=[1,1,0,1,1,0,0,1,0,0]$ (Fig. 3(a))
$p_{i}^{\prime}$ The DFS index of the inward edge, among all the inward edges, that is at the $i$-th position of $E(T)$. $p^{\prime}=[1,2,0,3,4,0,0,5,0,0]$ (Fig. 3(b))
$p^{\prime\prime}_{i}$ Replacing $p^{\prime}_{i}=0$ with $2n=10$ in $p^{\prime}$ to ignore the root case. $p^{\prime\prime}=$ $[1,2,10,3,4,10,10,5,10,10]$
$q_{ji}$ A binary variable which identifies the position $i$ of the non-zero input $x_{j}$ in $E(T)$, i.e., $q_{ji}=1$ if and only if $p^{\prime\prime}_{i}=x_{j}$. $q_{1,1}=q_{2,4}=1$, other variables are zero (Fig. 3(c))
$r^{\prime}_{ji}$ The required label of the inward edge of $x_{j}$. $r^{\prime}_{1,1}=3,r^{\prime}_{2,4}=2$, other variables are zero (Fig. 3(c))
Figure 3: Illustrations of the variables used Eqs.(1)-(5) in Lemma 1: (a) The variable $p_{i}$ which is 1 for the inward (black) edges and 0 for the outward (gray) edges in the directed tree corresponding to the tree $T$ given in Fig. 2(a), e.g., $p_{2}=1$ (resp., $p_{9}=0$) as there is an inward edge (resp., outward edge) with the DFS index 2 (resp., 9); (b) The variable $p^{\prime}_{i}$ (green), e.g., $p^{\prime}_{8}=5$ means that the inward edge of $5$ has the DFS index 8 in (a); (c) For a fixed DFS index $i$, the variable $q_{ji}=1$ for some $j$ (black inward edge), and $q_{ji}=0$ for all $j$ (gray inward edges), e.g., for the DFS index $i=4$, we have $j=2$ such that $q_{2,4}=1$ and thus the inward edge with the DFS index $4$ is depicted in black, whereas for $i=8$ there does not exist any $j$ such that $q_{j8}=1$, and so the inward edge with the DFS index $8$ is depicted in gray. The dark edges are the desired inward edges of which labels are required. The labels of these edges are stored by the variable $r^{\prime}_{ji}$, e.g., $r^{\prime}_{2,4}=2$ means that the desired inward edge specified by $x_{2}$ has the DFS index 44 and label 22.
Proposition 1 gives a necessary and sufficient condition for the $i$-th entry to be the outward edge of an inward edge at $\ell$-th position in an Euler string. The condition essentially depends on the number of inward edges and outward edges between the two given positions $i$ and $\ell$. Before going into the details, for any two positions i{i} and $\ell$}, with $1\leq\ell<i\leq 2n$, we denote by $\rm{in}_{\ell i}$ (resp., $\rm{out}_{\ell i}$), the number of inward edges (resp., outward edges) $t_{k}$, $\ell<k<i$. Consider the tree $T$ given in Fig. 2(a) and its Euler string $E(T)$ given in Fig. 2(c), for $i=7$, $\ell=2$, $\rm{in}_{2,7}=2$ as there are two inward edges $t_{4},t_{5}$ and $\rm{out}_{2,7}=2$ as there are two outward edges $t_{3},t_{6}$ (see Fig. 4(a)). So in this case $\rm{in}_{2,7}=\rm{out}_{2,7}$. Similarly, for $i=9,\ell=3$, $\rm{in}_{3,9}=3>2=\rm{out}_{3,9}$ (see Fig. 4(b)) and for $i=9,\ell=5$, $\rm{in}_{5,9}=1<2=\rm{out}_{5,9}$ (see Fig. 4(c)).
Figure 4: An illustration of the number of inward edges (green) (resp., outward edges (blue)) for the edges $t_{i},i=7,9$ and $t_{\ell},\ell=2,3,5$ which are depicted in gray boxes.
Proposition 1.
Let $E(T)=t_{1},t_{2},\ldots,t_{2n}$ denote the Euler string of a tree $T$ with $n$ edges and $t_{i}\in\Sigma=\{1,2,\ldots,m\}$. Then $t_{i}$, $1\leq i\leq 2n$, is an outward edge of $t_{\ell}$ if and only if (i)-(iv) hold
  1. (i)
    ${\ell}\in[1,i-1]$,
  2. (ii)
    $t_{i}=t_{\ell}+m$,
  3. (iii)
    $\rm{in}_{\ell i}=\rm{out}_{\ell i}$, and
  4. (iv)
    $\ell$} is the largest index that satisfies (i)-(iii).
Proof.
We know that the Euler string follows the DFS. Therefore the DFS index of all descendant edges of a given inward edge appear between the inward edge and its outward edge from which the result follows. ∎
A demonstration of Proposition 1 is given in Example 2.
Example 2.
Consider the tree $T$ given in Fig. 2(a) and its Euler string $E(T)$ shown in Fig. 2(c). We want to determine if the edges $t_{i}$, $i=4,7,9$ in $E(T)$ are outward edges of some inward edges by using Proposition 1. We see that Proposition 1(ii) does not hold for $i=4$ and any $\ell\in[1,3]$, therefore $t_{4}$ is not an outward edge of any inward edge, which is consistent with the fact. For $i=7$, Proposition 1(ii) and (iii) are satisfied for both $\ell=2,4$. For $\ell=2$ (resp., $\ell=4$), $\rm{in}_{2,7}=2=\rm{out}_{2,7}$ (resp., $\rm{in}_{4,7}=1=\rm{out}_{4,7}$) (see Fig. 4(a)). That is $t_{2}$ and $t_{4}$ are both candidate inward edges of $t_{7}$. By using Proposition 1(iv), $t_{7}$ is the outward edge of $t_{4}$. For $i=9$, we see that $t_{5}$ and $t_{8}$ are the only edges that satisfy Proposition 1(i)-(ii), but $t_{5}$ does not satisfy the Proposition 1(iii). Therefore $t_{8}$ is the inward edge of $t_{9}$ implying that $t_{9}$ is an outward edge.
The existence of a ReLU network to identify the positions and labels of outward edges in an Euler string based on Proposition 1 is discussed in Lemma 2.
Lemma 2.
Let $T$ be a tree with $n$ edges, and $x=x_{1},x_{2},\ldots,x_{d}$ be a random sequence of integers over the interval $[0,n]$. Then there exists a ReLU network with size $\mathcal{O}(dn^{2})$ and constant depth that can identify the position and label of outward edge of each non-zero input $x_{j}$ in the Euler string $E(T)$.
Proof.
Let $E(T)=t_{1},t_{2},\ldots,t_{2n}$. The proof completes by expressing the conditions of Proposition 1 in terms of ReLU activation function. Proposition 1 requires the labels of the inward edges of each $x_{j}\neq 0$ which can be computed by using Eqs.(1)-(5) of Lemma 1. Then the positions and labels of the required outward edges in $E(T)$ can be obtained by using the following system of equations (see Example 3), where $i,\ell\in\{1,2,\ldots,2n\}$, $j\in\{1,2,\ldots,d\}$ and $C$ is a large number.
$r_{i}$ $=t_{i}\cdot p_{i},$ (6)
$s_{i}$ $=t_{i}-r_{i},$ (7)
$v_{{\ell}i}$ $=\begin{cases}0&\text{~if~}i\leq{\ell},\\ \max(\delta(s_{i},r_{\ell}+m)-\\ ~~C(\sum_{k={\ell}+1}^{i-1}H(s_{k}-1)-\\ ~~\sum_{k={\ell}+1}^{i-1}H(r_{k}-1)),0)&\text{otherwise},\end{cases}$ (8)
$v^{\prime}_{{\ell}i}$ $=\delta(v_{{\ell}i},1),$ (9)
$w_{{\ell}i}$ $=\max(v^{\prime}_{{\ell}i}-\sum_{k={\ell}+1}^{i-1}v^{\prime}_{ki},0),$ (10)
[…]
$z_{j}$ $=i\cdot\sum_{i=1}^{2n}\sum_{{\ell}=1}^{2n}w^{\prime}_{j{\ell}i},$ (12)
$z^{\prime}_{ji}$ $=t_{i}\cdot\sum_{{\ell}=1}^{2n}w^{\prime}_{j{\ell}i}.$ (13)
The non-zero variable $r_{i}$ (resp., $s_{i}$) in Eq. (6) (resp., Eq. (7)) stores the label of the inward edge (resp., outward edge) at the $i$-th position of $E(T)$. Eqs. (8) and (9) encode Proposition 1(ii) and (iii), and Eq. (10) encodes Proposition 1(iv). In Eq. (11), $w^{\prime}_{j\ell i}=1$ if and only if $s_{i}=r^{\prime}_{j\ell}+m$ and $r^{\prime}_{j\ell}$ is the largest index with this property, i.e., all conditions of Proposition 1 are satisfied. Eq. (12) (resp., Eq. (13)) determines the positions (resp., labels) in $E(T)$ of the desired outward edges. These equations involve the maximum function, $\delta$ function and Heaviside function which can be simulated by the ReLU activation function based on Proposition 1 by Ghafoor and Akutsu [33] and Theorem 1 by Kumano and Akutsu [23]. Therefore we can construct a twelve-layer neural network with ReLU as an activation function with size $\mathcal{O}(dn^{2})$ and constant depth that can identify the positions and labels of the desired outward edges of non-zero $x_{j}$. ∎
A demonstration of Lemma 2 is given in Example 3.
Example 3.
Reconsider the tree $T$ shown in Fig. (2)(a) with $E(T)=3,2,7,2,$ $4,9,7,4,9,8$, and the random sequence $x=1,3,0$. We wish to identify the positions and labels of the outward edges of $x_{j}\neq 0$ in $E(T)$. Note that the positions of outward edges are their DFS indices in the directed tree corresponding to $T$ as demonstrated in Figs. (2)(b) and (c). For $x_{1}=1$ (resp., $x_{2}=3$), the positions and labels of the outward edges are $10$ and $8$ (resp., $7$ and $7$), resp., whereas $x_{3}=0$ does not correspond to any outward edge, and therefore it is ignored. We discuss the variables used in Lemma 2 to get the required positions and labels as follows. The variables $p_{i},p_{i}^{\prime},p_{i}^{\prime\prime},q_{ji},r_{ji}^{\prime}$ used in Eqs. (1)-(5) are discussed, in detail, in Example 1 and Fig. 3. We discuss the variables of Eqs. (6)-(13). An illustration of these variables is given in Fig. (5). In the rest of the discussion, more than one subscripts are separated by the commas to avoid confusion.
Figure 5: An illustration of the variables used in Eqs. (6)-(13) of Lemma 2 to identify the positions and labels of the desired outward edges.
$x_{j}$ Specify the outward edge of $x_{j}$ of which the position and label are required. In this case $x=1,3,0$, which is illustrated in Fig. 2(a).
$p_{i},p_{i}^{\prime},p_{i}^{\prime\prime},q_{ji},r_{ji}^{\prime}$ are explained in Example 1, Table 2 and Fig. 3.
$r_{i}$ The label of the inward edge of $i$, if it exists, e.g., in Fig. 5(a), $r_{1}=3$ (resp., $r_{10}=0$) as there exists (resp., does not exist) an inward edge of $1$ (resp., $10$). The label of the inward edge of $1$ is $3$. The values of these variables are listed in $r=[3,2,0,2,4,0,0,4,0,0]$.
$s_{i}$ The label of the outward edge of $i$, if it exists, e.g., in Fig. 5(a), $s_{10}=8$ (resp., $s_{1}=0$) as there exists (resp., does not exist) an outward edge of $10$ (resp., $1$). The label of the outward edge of $10$ is $8$. Similarly, we get $s=[0,0,7,0,0,9,7,0,9,8]$.
$v_{\ell i}$ A variable which can take a value from $\{1,C,0\}$: $v_{\ell i}=1$ if the conditions of Proposition 1(ii) and (iii) are satisfied, e.g., $i=7,\ell=2,4$, $\rm{in}_{\ell i}=\rm{out}_{\ell i}$ (see Example 2 and Fig. 4(a)); $v_{\ell i}=C$ if Proposition 1(iii) is violated with the number of number of inward edges greater than the number of outward edges, i.e., $\rm{in}_{\ell i}>\rm{out}_{\ell i}$ holds, e.g., $v_{3,9}=C$ as $\rm{in}_{3,9}>\rm{out}_{3,9}$ (see Fig. 4(b)); $v_{\ell i}=0$ if Proposition 1(iii) is violated with the number of inward edges less than that of outward edges, i.e., $\rm{in}_{\ell i}<\rm{out}_{\ell i}$ holds, e.g., $v_{5,9}=0$ as $\rm{in}_{5,9}<\rm{out}_{5,9}$ (see Fig. 4(c)). The values $v_{2,7}=v_{4,7}=1$, $v_{5,7}=0$ and $v_{1,7}=C$ are illustrated in Fig. 5(b). Thus we have, $v_{1,10}=v_{2,3}=v_{2,7}=v_{4,7}=v_{5,6}=v_{8,9}=1$, $v_{1,3}=v_{1,5}=v_{1,6}=v_{1,7}=v_{1,9}=v_{2,6}=v_{3,5}=v_{3,6}=v_{3,7}=v_{3,9}=v_{4,6}=v_{7,9}=C$, and other variables are zero.
$v^{\prime}_{\ell i}$ Replacing $C$ with 0 in $v_{\ell i}$, e.g., $v_{2,6}=C$, and therefore $v^{\prime}_{2,6}=0$ (see Fig. 5(b)). Thus $v^{\prime}_{1,10}=v^{\prime}_{2,3}=v^{\prime}_{2,7}=v^{\prime}_{4,7}=v^{\prime}_{5,6}=v^{\prime}_{8,9}=1$, and other variables are zero.
$w_{\ell i}$ A binary variable which is one when $\ell$ is the largest number that satisfies Proposition 1(i)-(iii), i.e., Proposition 1(iv)is satisfied, e.g., $v^{\prime}_{2,7}=v^{\prime}_{4,7}=1$ are the only non-zero variables for $i=7$. Since $4$ is largest among such variables, we have $w_{4,7}=1$ (see Fig. 5(b)). Similarly, $w_{1,10}=w_{2,3}=w_{4,7}=w_{5,6}=w_{8,9}=1$, and other variables are zero.
$w^{\prime}_{j\ell i}$ A binary variable to identify the desired outward edges corresponding to $x_{j}\neq 0$. More precisely $w^{\prime}_{j\ell i}$ is one when $t_{i}$ is the outward edge of the inward edge $t_{\ell}$ (Proposition 1 is satisfied), and $t_{\ell}$ is the inward edge of $x_{j}$, e.g., $w_{2,4,7}=1$ because $t_{7}$ is the outward edge of the inward edge $t_{4}$ which corresponds to $x_{2}=3$ (see Fig. 3(c)). In Fig. 5(c) the outward edges that have non-zero (resp., zero) value of $w^{\prime}_{j\ell i}$ are depicted by dark (resp., gray) edges. Thus, $w^{\prime}_{1,1,10}=w^{\prime}_{2,4,7}=1$, and other variables are zero.
$z_{j}$ Position $i$ of the outward edge of $x_{j}\neq 0$. In Fig. 5(c), $z_{1}=10$ and $z_{2}=7$ means that the position of the outward edges of $x_{1}$ (resp., $x_{2}$) is $10$ (resp., $7$). Thus $z=[10,7,0]$.
$z^{\prime}_{ji}$ Label of the outward edge of $x_{j}\neq 0$ which is at the position $i$. In this case $z^{\prime}_{1,10}=8,z^{\prime}_{2,7}=7$, and other variables are zero (see Fig. 5(c)).
4 TSd-generative ReLU
Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$, and a non-negative integer $d$. We define the TSd-generative ReLU to be a ReLU neural network with $2d$ input nodes $x=x_{1},x_{2},\ldots,x_{2d}$ over $\{0,\ldots,n\}$, and $2n$ output nodes $u=u_{1},u_{2},\ldots,u_{2n}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance at most $2d$ from $E(T)$ can be obtained by the substitution of appropriate $x_{j+d}$ and $x_{j+d}+m$ at the inward edge and outward edge of $x_{j}\neq 0$, respectively. In this context $x_{1},\ldots,x_{d}$ represents the inward edges for the substitution operations, while $x_{d+1},\ldots,x_{2d}$ denotes the values to be substituted during these operations. An illustration of such a network is given in Fig. 6, where $m=5$, $d=3$, $x=1,3,0,5,1,2$, means that the labels of the inward edges (resp., outward edges) of $x_{1}=1,x_{2}=3$ are substituted by $x_{4}=5,x_{5}=1$ (resp., $10,6$). The entry $x_{3}=0$ corresponds to the root, and is ignored.
Figure 6: An illustration of a TSd-generative ReLU with the input layer $x=1,3,0,5,1,2$, $E(T)=3,2,7,2,4,9,7,4,9,8$ and output layer $u=E(U)=5,2,7,1,4,9,6,4,9,10$. The substitution operations on $E(T)$ and $E(U)$ are depicted with green boxes and red values, respectively.
The existence of TSd-generative ReLU network is discussed in Theorem 1.
Theorem 1.
For a rooted ordered tree $T$ of size $n+1$ with a label set $\Sigma=\{1,2,\ldots,m\}$, and a non-negative integer $d$, there exists a TSd-generative ReLU network with size $\mathcal{O}(dn^{2})$ and constant depth.
A proof and an explanation of each variable of Theorem 1 is given in Example 4 in Appendix 9.
5 TDd, TId-generative ReLU
Let $T$ be a tree with $n+1$ nodes, labels from $\Sigma=\{1,2,\ldots,m\}$ and Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$, and a non-negative integer $d$. We define the TDd-generative ReLU to be a ReLU neural network that can generate all Euler strings over $\Sigma$ with the tree edit distance at most $2d$ from $E(T)$ by deleting from $E(T)$, the appropriate inward and outward edges of $x=x_{1},x_{2},\ldots,x_{d}$, over $\{0,\ldots,n\}$. To output a fixed number of nodes, we assume that $E(T)$ is padded with $2d$ $B$s, $B\gg m$, so that the network can delete the inward and outward edges of $x_{j}\neq 0$ and delete two $B$s corresponding to each $x_{j}=0$. We call such a string a padded Euler string. In this setting, we fix $2n$ output nodes $y=y_{1},\ldots,y_{2n}$ of the network, where $y_{1},\ldots,y_{2n-2d^{\prime}}$, $d^{\prime}\leq d$ is the Euler string of some tree $U$ with the tree edit distance $2d$ from $E(T)$ by deleting $d^{\prime}$ inward and outward edges of non-zero entries of $x$, and the remaining $2d^{\prime}$ entries $y_{2n-2d^{\prime}+1},\ldots,y_{2n}$ are $B$s. An illustration of a TDd-generative ReLU is given in Fig. 7, where $m=5$, $d=3$, $x=1,3,0$, and $T$ is given in Fig. 2 with $E(T)=3,2,7,2,4,9,7,4,9,8$, and the padded $E(T)$ with six $B$s. The network will delete the inward and outward edges of $1$ and $3$ as depicted in the figure, and delete two $B$s corresponding to 0. The resultant string is $y=2,7,4,9,4,9,B,B,B,B$, and by removing $B$s from $y$ we get the desired Euler string $E(U)=2,7,4,9,4,9$ of the tree $U$ as shown in the Fig. 7.
Figure 7: An illustration of a TDd-generative ReLU with the input layer $x=1,3,0$, padded Euler string $3,2,7,2,4,9,7,4,9,8,B,B,B,B,B,B$ and the output layer $2,7,4,9,4,9,B,B,B,B$. By deleting $B$s we can get the resultant string $E(U)=2,7,4,9,4,9$ obtained by deleting the inward and outward edges of $x=1,3$.
The existence of TDd-generative ReLU network is discussed in Theorem 2.
Theorem 2.
For a rooted ordered tree $T$ of size $n+1$ with nodes from $\Sigma=\{1,2,\ldots,m\}$, and a non-negative integer $d$, there exists a TDd-generative ReLU network with size $\mathcal{O}(n^{2})$ and constant depth.
A proof and an explanation of each variable of Theorem 2 is given in Example 5 in Appendix 9.
We define TId-generative ReLU as follows. Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$. We define the TId-generative ReLU to be a ReLU neural network with $4d$ input nodes $x=x_{1},x_{2},\ldots,x_{4d}$ over $\{0,\ldots,n\}$, and $2n+2d$ output nodes $u=u_{1},u_{2},\ldots,u_{2n+2d}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance exactly $2d$ from $E(T)$ can be obtained by the insertion of appropriate child nodes $x^{\prime}_{j}$ of $x_{j}$ with inward edges and outward edges of labels $x_{j+3d}$ and $x_{j+3d}+m$, resp., of $x^{\prime}_{j}$, and setting the appropriate ($x_{j+d}$, $x_{j+d}+1$, …, $x_{j+2d}$)-th children of the node $x_{j}$ as the children of $x^{\prime}_{j}$. In this study, we do not consider inserting new nodes as children of a newly inserted node. In this context, $x_{1},\ldots,x_{d}$ represent the nodes for the insertion operations, $x_{1+d},\ldots,x_{d+d}$ and $x_{1+2d},\ldots,x_{d+2d}$ represent the lower and upper bounds to determine the subsequences of children that will be set as the children of the inserted nodes, and $x_{1+3d},\ldots,x_{d+3d}$ represents the labels to be inserted. For convenience, we denote $x_{1},\ldots,x_{d}$ , $x_{1+d},\ldots,x_{2d}$, $x_{1+2d},\ldots,x_{3d}$, and $x_{1+3d},\ldots,x_{4d}$ by $x^{1}$, $x^{2}$, $x^{3}$, and $x^{4}$, respectively. Note that some lower and upper bounds of children may not be valid due to the random nature of $x$, and thus need to be refined to perform appropriate insertion operations. Such invalid bounds, their refinements and appropriate insertions are listed in Table 3, where $D(x_{j}^{1})$ denotes the number of children of the node $x_{j}^{1}$.
Figure 8: An illustration of a TId-generative ReLU with the input layer $x=1,0,3,0,2,4,1,1,3,2,5,1,4,1,3,5$, $E(T)=3,2,7,2,4,9,7,4,9,8$ and output layer u=E(U)=1,6,5,3,2,7,4,2,4,9,3,8,7,4,9,9,8,10u=E(U)=1,6,5,3,2,7,4,2,4,9,3,8,7,4,9,9,8,10. The insertions are depicted in red in $U$.
Table 3: Invalid bounds of children, their refinements and appropriate insertions.
S. no. Invalid bounds Refinements Insertions
(i) $D(x_{j}^{1})<x_{j}^{2}$ $x_{j}^{2}:=0$ Insert a leaf before the first child of $x_{j}^{1}$
(ii) $D(x_{j}^{1})<x_{j}^{3}$ $x_{j}^{3}:=0$ Insert a leaf after the $x_{j}^{2}$-th child of $x_{j}^{1}$
(iii) $x_{j}^{2}>x_{j}^{3}$ $x_{j}^{3}:=0$ Insert a leaf after the $x_{j}^{2}$-th child of $x_{j}^{1}$
(iv) $x_{j}^{1}=x_{k}^{1},j<k,x_{j}^{3}>x_{k}^{2}$ $x_{j}^{3}:=0$ Insert a leaf after the $x_{j}^{2}$-th child of $x_{j}^{1}$
(v) $x_{j}^{1}=x_{k}^{1},j>k,x_{j}^{2}=x_{k}^{3}$ $x_{j}^{3}:=0$ Insert a leaf after the $x_{j}^{2}$-th child of $x_{j}^{1}$
(vi) $x_{j}^{1}=x_{k}^{1},j<k,x_{j}^{2}>x_{k}^{2}$ $x_{j}^{2}:=0$ Insert a leaf before the first child of $x_{j}^{1}$
(vii) $x_{j}^{1}=x_{k}^{1},x_{j}^{2}=x_{k}^{2},x_{j}^{3}>x_{j}^{2}$ $x_{j}^{2}:=0$ Insert two leaves before the first child of $x_{j}^{1}$
(viii) $x_{j}^{2}=0$ $x_{j}^{3}:=0$ Insert a leaf before the first child of $x_{j}^{1}$
(ix) $x_{j}^{2}\neq 0,x_{j}^{3}=0$ $x_{j}^{2}:=x_{j}^{2}+1$ Insert a leaf after the $x_{j}^{2}$-th child of $x_{j}^{1}$
An illustration of a TId-generative ReLU is given in Fig. 8, where $m=5$, $d=4$ and x=x=1, 0, 3, 0, 2, 4, 1, 1, 3, 2, 5, 1, 4, 1, 3, 5. For convenience, we perform the insertion operations in the ascending order of the values $x^{1}$, i.e., in this case, the insertion is performed by considering the sequence $0,0,1,3,4,1,2,1,2,1,3,5,1,5,4,3$. We discuss the insertion process as follows. For the node $x_{2}^{1}=0$, the bounds are $x_{2}^{2}=4$ and $x_{2}^{3}=2$, which are invalid as $D(x_{2}^{1})=1$. Therefore by applying refinements (i) and (ii), we set $x_{2}^{2}:=0$ and $x_{2}^{3}:=0$, and thus insert a leaf with label $1$ and DFS index $1$ (see Fig. 8), inward and outward edges with labels 1 and 6 at 1st and 2nd positions of the resultant Euler string $E(U)$, respectively. The bounds for the node $x_{4}^{1}=0$ are valid, and hence a new node with label $5$ is inserted with index $2$, as shown in Fig. 8, and insert inward and outward edges 5 and 10 at 3rd and 18th positions of $E(U)$. For the node $x_{1}^{1}=1$, the given lower bound and upper bound for the children are $x_{1}^{2}=2$ and $x_{1}^{3}=3$, which are valid as the number $D(x_{1}^{1})$ of children of $x_{1}^{1}$ are 3 as depicted in Fig. 8. The 2nd and 3rd children of $x_{1}^{1}$ have the DFS indices $3$ and $5$, respectively. Thus a new node ${x^{\prime}}_{1}^{1}$ with label $4$ and index $5$ is inserted by setting the 2nd and 3rd children of $x_{1}^{1}$ as the children of ${x^{\prime}}_{1}^{1}$. The revised indices of the children are $6$ and $9$, resp., as shown in Fig. 8. Inward and outward edges with labels 4 and 9 are inserted at 7th and 16th position of $E(U)$, respectively. For the node $x_{3}^{1}=3$, the bounds are $x_{3}^{2}=1$ and $x_{3}^{3}=5$, where the upper bound is invalid as $D(x_{3}^{1})=1$. By applying the refinement (iii), we set $x_{3}^{3}:=0$, and by (ix) we have $x_{3}^{2}:=2$, therefore insert a leaf after the first child of $x_{3}^{1}$ with DFS index $8$, and insert inward and outward edges with labels 3 and 8 at the 11th and 12th positions of $E(U)$, respectively. The resultant tree $U$ has the Euler string $E(U)=1,6,5,3,2,7,4$, $2,4,9,3,8,7,4,9,9,8,10$.
The existence of TId-generative ReLU network is discussed in Theorem 3.
Theorem 3.
For a rooted ordered tree $T$ of size $n+1$ with nodes from $\Sigma=\{1,2,\ldots,m\}$, and a non-negative integer $d$, there exists a TId-generative ReLU network with size $\mathcal{O}(n^{3})$ and constant depth.
A proof of Theorem 3 and an explanation of each variable used in it are given in Example 6 in Appendix 9.
6 TEd-generative ReLU
Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$. We define the TEd-generative ReLU to be a ReLU neural network such that each Euler string over $\Sigma$ with edit distance at most $2d$ from $E(T)$ due to deletion, substitution and insertion operations can be obtained by appropriately choosing an input $x=x_{1},x_{2},\ldots,x_{7d}$ of $7d$ nodes with $x_{j}\in[0,1)$, where $x_{j}$ is of the form $\cdot\Delta, ii$ is an integer and $\Delta$ is a small constant. The input $x_{1},x_{2},\ldots,x_{d}$ (resp., $x_{d+1},x_{d+2},\ldots,x_{3d}$ and $x_{3d+1},x_{3d+2},\ldots,x_{7d}$) represents the nodes for deletion (resp., substitution and insertion) operations. As a preprocessing step, the random inputs $x_{j}$ for $1\leq j\leq 2d$ and $3d+1\leq j\leq 6d$ (resp., $2d+1\leq j\leq 3d$ and $6d+1\leq j\leq 7d$) are converted into integers $i\in\{0,\ldots,n\}$ (resp., $\ell\in\Sigma$) if $x_{j}\in((i-1)/n,i/n]$ (resp., $x_{\ell}\in[(\ell-1)/m,\ell/m]$ for $\ell=1$, $((\ell-1)/m,\ell/m]$ otherwise). For example, when $n=5$ and $m=10$, the conversion table is given in Table 7 in Appendix 9. To output a fixed number of nodes, we assume that $E(T)$ is padded with $2d$ $B$s, where $B\gg\max(m,n)$. The network outputs $2n+2d$ nodes $y=y_{1},\ldots,y_{2n+2d}$ from which the desired string $E(U)$ can be obtained by trimming all $B$s from the start and end. More precisely, if $d_{1}$ (resp., $d_{2}$) deletion (resp., insertion) operations are performed, then $2d-2d_{1}$ (resp., $2d_{2}$}) number of $B$s will be trimmed from the end (resp., start) of the output $y$ as shown in Fig. 9 in Appendix 9.
Figure 9: An illustration of a TEd-generative ReLU for $d=3$, with the input layer x=0.3,0,0.38,0,0.46,0.55,0,0.6,0.88,0.66,0.75,0,0.55,0.87,0.03,0.02x=0.3,0,0.38,0,0.46,0.55,0,0.6,0.88,0.66,0.75,0,0.55,0.87,0.03,0.02, $0.45,0.09,0,0.7,0.5$, padded Euler string $3,2,12,2,4,14,12,4,14,13,B,B,B,B,B,B$ and the output layer $B,B,B,B,5,3,2,6,16,12,4,14,13,15,B,B,B,B$ with $d_{1}=d_{2}=1$. By trimming $B$s, we can get the resultant string $E(U)=5,3,2,6,16,12,4,14,13,15$ obtained by deleting, substituting and inserting the indicated nodes.
The existence of TEd-generative ReLU network is discussed in Theorem 4.
Theorem 4.
For a rooted ordered tree $T$ of size $n+1$ with nodes from $\Sigma=\{1,2,\ldots,m\}$, and a non-negative integer $d$, there exists a TEd-generative ReLU network with size $\mathcal{O}(n^{3})$ and constant depth.
A proof of Theorem 4, and an explanation of each variable used in Theorem 4 is given in Example 7 in Appendix 9.
7 Computational Experiments
We implemented the proposed networks on a machine with an AMD Ryzen 7 4800H, Radeon Graphics processor (2.90 GHz), 16 GB of RAM, and Windows 11 Pro using Python version 3.11.6.
Figure 10: Trees used in the computational experiments of TId-generative ReLU and TEd-generative ReLU.
For each tree $T_{i}$, $i=1,2,3,4$ illustrated in Figs. 10(a)-(d) with $8,10,6,11$ nodes, labels from $\Sigma=\{1,2,\ldots,10\}$, and $d=2,2,3,2$, resp., we generated all non-isomorphic trees $U$ that have tree edit distance exactly $d$ by using the proposed TId-generative ReLU networks. The newly inserted nodes in $T_{1}$ are 7, 7, in $T_{2}$ are 9, 9, in $T_{3}$ are 8, 8, 8, and in $T_{4}$ are 2, 2. The computational results such as the number of nodes in each layer of the constructed TId network along with the number of generated trees for each input tree by the network are provided in Table 4. A summary of these computational results is given below. For $T_{i}$, let $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm I}_{i}$ denote the sequence of number of hidden layers, total number of hidden nodes, minimum number of hidden nodes, average number of hidden nodes, and maximum number of hidden nodes, resp., in the TId-generative ReLU network for $T_{i}$. From these experiments, we have $(57,32876,6,576.77,11903)^{\rm I}_{1}$, $(57,57324,6,1005.68,24841)^{\rm I}_{2}$, $(57,20556,9,360.63,4500)^{\rm I}_{3}$, $(57,73064,6,1281.82,33860)^{\rm I}_{4}$. These values are illustrated in Fig. 11(a) for each tree $T_{i}$, $i=1,2,3,4$. Observe that the depth remains fixed across all trees, confirming Theorem 3. Moreover, the widest layers are significantly larger than the others; for example, in $T_{4}$ with 11 nodes, the widest layer contains 33860 nodes which is over 26.41 times the average layer size (1281.82). Additionally, the total number of nodes grows faster than the tree size, e.g., while 11 nodes are roughly twice 6, the corresponding total node count (73064) is 3.56 times that for 6 nodes (20556).
Figure 11: Information of the architectures of the proposed ReLU-networks for insertion and unified cases and their comparison: (a) Information of $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm I}_{i}$ of the proposed ReLU-networks for insertion case for each tree $T_{i}$, $i=1,2,3,4$; (b) Information of $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm E}_{i}$ of the proposed ReLU-networks for unified case for each tree $T_{i}$, $i=1,2,3,4,5$; (c) A comparison of the information of $(\#\rm{TN},\rm{AvgN},\#\rm{MaxN})_{i}$ of the proposed ReLU-networks for insertion and unified cases for each tree $T_{i}$, $i=1,2,3,4$.
Similarly, TEd-generative ReLU neural networks were constructed for the trees $T_{i}$, $i=1,2,3,4,5$ given in Figs. 10(a)-(e), where $T_{5}$ has $21$ nodes, labels from $\Sigma=\{1,2,\ldots,10\}$, and $d=2$. The newly inserted (resp., substituted) nodes in $T_{1}$ are 0.7, 0.7 (resp., 0.55, 0.55), in $T_{2}$ are 0.9, 0.9 (resp., 0.7, 0.7), in $T_{3}$ are 0.8, 0.8, 0.8 (resp., 0.88, 0.88, 0.88), in $T_{4}$ are 0.2, 0.2 (resp., 0.9, 0.9) and in $T_{5}$ are 0.3, 0.3 (resp., 0.99, 0.99).
Table 4: Experimental results for TId-generative ReLU
Input trees Size of each hidden layer of TId-generative ReLU Number of generated trees
$T_{1}$, Fig. 10(a) 36, 106, 806, 666, 232, 902, 400, 204, 23, 848, 218, 428, 218, 1073, 443, 758, 11903, 458, 670, 377, 252, 524, 520, 528, 524, 520, 544, 528, 520, 540, 528, 518, 516, 528, 518, 518, 516, 640, 962, 482, 6, 26, 10, 22, 12, 6, 278, 23, 22, 352, 82, 90, 50, 26, 262, 78, 36 318
$T_{2}$, Fig. 10(b) 44, 134, 1322, 1070, 368, 1446, 656, 332, 27, 1376, 350, 692, 350, 1737, 711, 1186, 24841, 730, 1074, 583, 392, 812, 808, 816, 812, 808, 832, 816, 808, 828, 816, 806, 804, 816, 806, 806, 804, 960, 1522, 762, 6, 26, 10, 22, 12, 6, 342, 27, 26, 432, 102, 94, 54, 30, 330, 98, 44 518
$T_{3}$, Fig. 10(c) 32, 82, 422, 362, 132, 529, 212, 112, 23, 452, 122, 232, 122, 573, 243, 474, 4500, 320, 378, 258, 164, 453, 447, 465, 456, 447, 501, 465, 447, 489, 459, 441, 438, 456, 441, 441, 438, 576, 795, 399, 9, 51, 15, 45, 27, 9, 345, 23, 22, 436, 88, 172, 88, 28, 286, 82, 32 546
$T_{4}$, Fig. 10(d) 48, 148, 1628, 1308, 448, 1766, 808, 408, 29, 1688, 428, 848, 428, 2129, 869, 1436, 33860, 890, 1312, 704, 474, 980, 976, 984, 980, 976, 1000, 984, 976, 996, 984, 974, 972, 984, 974, 974, 972, 1144, 1850, 926, 6, 26, 10, 22, 12, 6, 374, 29, 28, 472, 112, 96, 56, 32, 364, 108, 48 660
The computational results of these experiments are given in Table 5 which are summarized below. For $T_{i}$, let $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm E}_{i}$ denote the sequence of number of hidden layers, total number of hidden nodes, minimum number of hidden nodes, average number of hidden nodes, and maximum number of hidden nodes, resp., in the TEd-generative ReLU network for $T_{i}$. For the computational results listed in Table 5 we have $(108,50360,14,466.30,11917)^{\rm E}_{1}$, $(108,82060,14,759.81$, $24859)^{\rm E}_{2}$, $(108,35803,19,331.51,4510)^{\rm E}_{3}$, $(108,101930,14$, $943.80,33880)^{\rm E}_{4}$, $(108,493790,14,4572.13,263350)^{\rm E}_{5}$ for TEd-generative ReLU networks. These values are illustrated in Fig. 11(b) for each tree $T_{i}$, $i=1,2,3,4,5$.
Moreover a comparison of the TId and TEd is given in Fig. 11(c). Observe that TEd networks are deeper and, on average, narrower than TId networks for the same tree. Similarly, TEd networks exhibit larger minimum widths compared to TId networks, while the maximum layer widths are nearly identical in both cases. These findings indicate that the insertion operation contributes the most to the overall size of the unified networks, which are composed of substitution, deletion, and insertion components. This observation also supports Theorems 3 and 4.
Table 5: Experimental results for TEd-generative ReLU
Input trees Size of each hidden layer of TEd-generative ReLU Number of generated trees
$T_{1}$, Fig. 10(a) 2016, 28, 120, 42, 34, 40, 42, 46, 30, 24, 22, 14, 66, 144, 1418, 482, 392, 1470, 696, 1650, 678, 354, 48, 30, 84, 30, 102, 30, 292, 82, 26, 70, 156, 54, 124, 908, 374, 262, 934, 458, 1816, 640, 444, 80, 164, 80, 276, 78, 36, 22, 50, 148, 50, 120, 890, 356, 218, 806, 414, 218, 37, 862, 232, 442, 232, 1087, 457, 772, 11917, 472, 684, 391, 266, 538, 534, 542, 538, 534, 558, 542, 534, 554, 542, 532, 542, 532, 532, 530, 654, 976, 496, 20, 40, 24, 36, 26, 20, 292, 37, 36, 366, 82, 90, 50, 26, 262, 78, 36 747
$T_{2}$, Fig. 10(b) 2240, 28, 120, 42, 34, 40, 42, 46, 30, 24, 22, 14, 74, 172, 2082, 674, 564, 2146, 1024, 2454, 1002, 518, 56, 34, 100, 34, 122, 34, 372, 102, 30, 82, 196, 66, 156, 1452, 550, 406, 1486, 730, 2980, 1036, 712, 100, 208, 100, 352, 98, 44, 26, 62, 188, 62, 152, 1430, 528, 350, 1322, 674, 350, 45, 1394, 368, 710, 368, 1755, 729, 1204, 24859, 748, 1092, 601, 410, 830, 826, 834, 830, 826, 850, 834, 826, 846, 834, 824, 834, 824, 824, 822, 978, 1540, 780, 24, 44, 28, 40, 30, 24, 360, 45, 44, 450, 102, 94, 54, 30, 330, 98, 44 1223
$T_{3}$, Fig. 10(c) 2688, 42, 204, 63, 51, 60, 63, 69, 45, 36, 33, 21, 89, 142, 1141, 405, 325, 1250, 562, 1314, 546, 290, 50, 34, 82, 34, 98, 34, 298, 88, 28, 84, 127, 48, 98, 498, 275, 165, 555, 265, 1355, 455, 355, 85, 175, 85, 295, 82, 32, 22, 42, 112, 42, 92, 482, 259, 122, 422, 222, 122, 33, 462, 132, 242, 132, 583, 253, 484, 4510, 330, 388, 268, 174, 463, 457, 475, 466, 457, 511, 475, 457, 499, 469, 451, 466, 451, 451, 448, 586, 805, 409, 19, 61, 25, 55, 37, 19, 355, 33, 32, 446, 88, 172, 88, 28, 286, 82, 32 $25$
$T_{4}$, Fig. 10(d) 2352, 28, 120, 42, 34, 40, 42, 46, 30, 24, 22, 14, 78, 186, 2462, 782, 662, 2532, 1212, 2916, 1188, 612, 60, 36, 108, 36, 132, 36, 412, 112, 32, 88, 216, 72, 172, 1772, 650, 490, 1810, 890, 3670, 1270, 870, 110, 230, 110, 390, 108, 48, 28, 68, 208, 68, 168, 1748, 626, 428, 1628, 828, 428, 49, 1708, 448, 868, 448, 2149, 889, 1456, 33880, 910, 1332, 724, 494, 1000, 996, 1004, 1000, 996, 1020, 1004, 996, 1016, 1004, 994, 1004, 994, 994, 992, 1164, 1870, 946, 26, 46, 30, 42, 32, 26, 394, 49, 48, 492, 112, 96, 56, 32, 364, 108, 48 1550
$T_{5}$, Fig. 10(e) 3472, 28, 120, 42, 34, 40, 42, 46, 30, 24, 22, 14, 118, 326, 8022, 2302, 2082, 8152, 3972, 9736, 3928, 1992, 100, 56, 188, 56, 232, 56, 812, 212, 52, 148, 416, 132, 332, 6732, 2090, 1770, 6810, 3370, 14530, 4930, 3330, 210, 450, 210, 770, 208, 88, 48, 128, 408, 128, 328, 6688, 2046, 1648, 6448, 3248, 1648, 89, 6608, 1688, 3328, 1688, 8289, 3369, 5296, 263350, 3410, 5052, 2614, 1774, 3580, 3576, 3584, 3580, 3576, 3600, 3584, 3576, 3596, 3584, 3574, 3584, 3574, 3574, 3572, 3904, 6930, 3486, 46, 66, 50, 62, 52, 46, 734, 89, 88, 912, 212, 116, 76, 52, 704, 208, 88 6309
For each tree $T_{i}$, the inputs $x$ and the corresponding Euler strings $E(U)$ generated by the TId and TEd networks are listed in the supplementary material S1 which is available on  https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
Additionally, we conducted experiments to generate trees with a given edit distance by using state-of-the-art graph generative models called GraphRNN by You et al. [53] and GraphGDP by Huang et al. [57] for comparison. For this purpose, we randomly generated datasets of sizes 100, 150, 150, 200, and 800 of trees whose distances from $T_{1}$, $T_{2}$, $T_{3}$, $T_{4}$ and $T_{5}$ are $2,2,3,2$ and $2$, respectively. For simplicity, the labels of the underlying trees were ignored. We trained GraphRNN (dependent Bernoulli variant) (resp., GraphGDP) on each of these datasets by using a 4-layer RNN (resp., 4-layer GNN) with hidden neuron size 128 (resp., 128). Training ran for around 3000 epochs with batch size 32, learning rate 0.003 for GraphRNN and 0.00002 for GraphGDP by using 80% of the input dataset, whereas 20% dataset was used for testing, and default parameters were retained for other settings. As a result, for each tree, $1024$ (resp., 512) samples were generated by using each GraphRNN (resp., GraphGDP). Computational results are given in Table 6.
The results demonstrate notable variability in the performance of both models across different tree structures. It is important to note that both GraphRNN and GraphGDP were expected to generate tree structures, yet neither model guarantees that all generated samples are valid trees. Additionally, among the generated trees, only those that satisfy the specified edit distance constraint are considered valid in this evaluation.
GraphRNN consistently generated a high percentage of trees (ranging from 86.3% to 96.6%), yet only a small fraction of these were valid with respect to the target distance constraint. For instance, for trees $T_{1}$, $T_{2}$, and $T_{4}$ with distance $d=2$, the proportion of valid trees remained low 6.4%, 6.9%, and 2.0%, respectively. In the case of $T_{5}$, GraphRNN failed to produce any valid tree, despite generating over 92% of trees. The only tree where GraphRNN achieved relatively high success was $T_{3}$ (with $d=3$), where 35.8% of the generated samples were valid.
In contrast, GraphGDP generated fewer trees overall, with percentages ranging between 3.7% and 48.6% across different trees. However, it was able to generate a higher proportion of valid trees among those it produced, especially in the case of $T_{3}$, where 47.8% of its outputs were valid. For the other trees, especially those with $d=2$, the valid tree percentages were noticeably lower: 9.2% for $T_{1}$, 3.9% for $T_{2}$, and 0.0% for both $T_{4}$ and $T_{5}$.
Table 6: Percentage of trees and valid trees generated by GraphRNN [53] and Huang et al. [57] with a given distance $d$
Tree $n+1$ $d$ GraphRNN [53] GraphGDP [57]
Trees Valid trees Trees Valid trees
$T_{1}$ 8 2 86.3% 6.4% 48.6% 9.2%
$T_{2}$ 10 2 96.6% 6.9% 35.0% 3.9%
$T_{3}$ 6 3 92.0% 35.8% 47.9% 47.8%
$T_{4}$ 11 2 94.1% 2.0% 32.8% 0.0%
$T_{5}$ 21 2 92.1% 0.0% 3.7% 0.0%
These findings suggest that GraphRNN and GraphGDP struggle to enforce the tree edit distance constraint, particularly as tree size and complexity increase, and may generate samples that are not even trees.
The datasets and the graphs generated by GraphRNN and GraphGDP are available in the supplementary materials S2 and S3, resp., which are available on https://github.com/MGANN-KU/TreeGen_ReLUNetworks
8 Conclusion
We study the existence of ReLU-based generative networks for producing trees similar to a given tree with respect to the tree edit distance. Our approach transforms a rooted, ordered, and vertex-labeled tree into a rooted, ordered, and edge-labeled directed tree. This directed tree is then encoded as an Euler string, which serves as both the input and output of the ReLU generative networks. First, we proved that there exists a ReLU network of size $\mathcal{O}(dn)$ and constant depth that can identify the labels of $d$ inward edges in the Euler string. Furthermore, we showed that the outward edges corresponding to these $d$ inward edges can be identified using a ReLU network of size $\mathcal{O}(dn^{2})$ and constant depth. Building on these results, we demonstrated that all similar trees generated through substitution (resp., deletion and insertion) operations can be constructed by ReLU networks of size $\mathcal{O}(dn^{2})$ (resp., $\mathcal{O}(n^{2})$ and $\mathcal{O}(n^{3})$), all with constant depth. Finally, we proved that there exists a ReLU network of size $\mathcal{O}(n^{3})$ and constant depth that can generate all trees with a distance at most $d$ to the original tree under combined substitution, deletion, and insertion operations. These findings provide a theoretical foundation towards construction of compact generative models and open new directions for efficient tree-structured data generation.
In this study, we do not consider scenarios where a newly inserted node becomes the parent of subsequent inserted nodes. This design choice simplifies the construction and supports tractable enumeration, but it limits the completeness of the editing model by excluding certain nested insertions. Addressing this limitation would extend the expressiveness of the framework and is a promising direction for future work. comparison with the state-of-the-art graph generative models GraphRNN by You et al. [53] and GraphGDP by Huang et al. [57] revealed that these models struggled to generate trees at a specified edit distance, particularly as tree size and structural complexity increased. For instance, GraphRNN and GraphGDP could not generate a single valid tree with 21 vertices and edit distance 2. While this experiment serves as an initial benchmark, it also underscores the need for further comparative evaluations with other models, such as TreeGAN and diffusion-based tree generators, to more clearly position the strengths of our proposed ReLU-based construction. On the other side, our construction demonstrates that our proposed ReLU network with constant depth and polynomial size can generate all trees within a given edit distance; however, the number of neurons in certain hidden layers increases rapidly with tree size. For instance, generating trees with 21 nodes can require layers containing over 263350 neurons. Although this wide structure ensures theoretical expressivity and completeness, it presents challenges for scalability, implementation, and deployment in resource-constrained environments. To address this limitation, future work may focus on strategies such as width pruning, parameter sharing, and compression to control and reduce the rapid growth in network width. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
Author contributions
Conceptualization, M.G. and T.A.; methodology, M.G. and T.A.; software, M.G. validation, T.A.; formal analysis, M.G. and T.A.; investigation, M.G. data curation, M.G.; writing—original draft preparation, M.G. and T.A.; writing—review and editing, M.G. and T.A..; supervision, T.A.; project administration, T.A. All authors have read and agreed to the published version of the manuscript.
Acknowledgments
The work of Tatsuya Akutsu was supported in part by Grants 22H00532 and 22K19830 from Japan Society for the Promotion of Science (JSPS), Japan. The authors would like to thank Dr. Naveed Ahmed Azam, Quaid-i-Azam University Pakistan, for the useful technical discussions.
Conflict of interest
All authors have no conflicts of interest in this paper.

### v3

@@
 In this paper, we theoretically establish the existence and construction of generative networks capable of producing trees similar to a given tree with respect to the tree edit distance.
-Specifically, for a given rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$, we prove that all rooted, ordered, and vertex-labeled trees over $\Sigma$ with tree edit distance at most $d$ from $T$ can be generated using a ReLU-based generative network with size $\mathcal{O}(n^{3})$ and constant depth.
+Specifically, for a given rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$, we prove that any rooted, ordered, and vertex-labeled tree over $\Sigma$ with tree edit distance at most $d$ from $T$ can be generated using an appropriate random input sequence to a ReLU-based generative network of size $\mathcal{O}(n^{3})$ and constant depth.
 The proposed networks were implemented and evaluated for generating trees with up to 21 nodes.
@@
 As a result, they cannot provide exact enumeration of the underlying combinatorial space, nor can they ensure complete validity or coverage of all possible structured instances.
-As a step towards addressing these issues we study the exact and deterministic generation of all rooted, ordered, and vertex-labeled trees similar to a given tree using neural networks and prove the existence of ReLU-based generative networks for tree edit distance.
-Given a rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from a symbol set $\Sigma$, we theoretically establish the existence of ReLU-activated generative networks capable of producing all rooted trees over $\Sigma$ within tree edit distance at most $d$ from $T$.
+As a step towards addressing these issues, we study the exact generation of rooted, ordered, and vertex-labeled trees similar to a given tree using neural networks, and prove the existence of ReLU-based generative networks for the tree edit distance problem.
+Given a rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from a symbol set $\Sigma$, we establish the existence of a ReLU-activated network that takes as input a finite random sequence $x$, and outputs a similar tree with edit distance at most $d$ from $T$.
+The network is deterministic in the sense that, for each random input sequence $x$, it systematically identifies and performs the specified substitution, deletion, and insertion operations to produce a unique output tree $T^{\prime}$ whose tree edit distance from $T$ is at most $d$ (see the proofs of Theorems 1, 2, 3, and 4 for details).
+Different random input sequences may lead to the same output tree.
+Nevertheless, by considering an appropriate collection of random input sequences that covers all admissible edit operations, every tree similar to $T$ within distance $d$ can be generated.
 The key idea of our approach is to first construct a directed, rooted, ordered, and edge-labeled tree based on $T$.
@@
 ReLU generative networks that can identify the indices and labels of the directed edges in the Euler string are discussed in Section 3.
-Existence of ReLU networks to generate all trees with tree edit distance at most $d$ due to substitution operations is discussed in Section 4.
-Existence of ReLU networks to generate all trees with edit distance at most (resp., exactly) $d$ due to deletion (resp., insertion) operations is discussed in Section 5.
-Generation of all trees with tree edit distance at most $d$ due to simultaneous application of deletion, substitution and insertion operations by using a ReLU network is discussed in Section 6.
+Existence of ReLU networks to generate any tree with tree edit distance at most $d$ due to substitution operations is discussed in Section 4.
+Existence of ReLU networks to generate any tree with edit distance at most (resp., exactly) $d$ due to deletion (resp., insertion) operations is discussed in Section 5.
+Generation of any tree with tree edit distance at most $d$ due to simultaneous application of deletion, substitution and insertion operations by using a ReLU network is discussed in Section 6.
 Computational experiments are discussed in Section 7.
@@
 Therefore in the rest of the paper, we will use a random sequence $x_{1},x_{2},\ldots,x_{d}$, $d\geq 1$, with integers $x_{j}\in[0,n]$, unless stated otherwise, to specify the DFS indices of the vertices under consideration in a tree with $n+1$ vertices.
-We ignore $x_{j}=0$ and repeated entries in the case of substitution and deletion operations.
+In this work, we focus on generation of similar trees with the same root as that of the input tree, therefore $x_{j}=0$ is ignored in substitution and deletion operations.
+Similarly, repeated entries in the case of substitution and deletion operations are ignored.
 An example random sequence is given in Fig.
@@
 3 Identification of Edge Labels by ReLU Network
-We focus on designing generative networks with ReLU as an activation function to generate all trees that are similar to a given tree. More precisely, we are interested in the following problem:
+We focus on designing generative networks with ReLU as an activation function to generate any tree that are similar to a given tree. More precisely, we are interested in the following problem:
 Input: A rooted, ordered, and vertex-labeled tree $T$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$.
-Output: Construct generative networks with ReLU as an activation function that can generate all rooted, ordered, and vertex-labeled trees over $\Sigma$ with tree edit distance at most $d$ from $T$.
+Output: Construct generative networks with ReLU as an activation function that can generate any rooted, ordered, and vertex-labeled tree over $\Sigma$ with tree edit distance at most $d$ from $T$.
 We target this problem by reducing the tree edit distance problem to the string edit distance problem by representing trees as their Euler strings as explained in Section 2.
@@
 The labels of these edges are stored by the variable $r^{\prime}_{ji}$, e.g., $r^{\prime}_{2,4}=2$ means that the desired inward edge specified by $x_{2}$ has the DFS index 44 and label 22.
-Proposition 1 gives a necessary and sufficient condition for the $i$-th entry to be the outward edge of an inward edge at $\ell$-th position in an Euler string.
+Proposition 1 gives a necessary and sufficient condition for the $i$-th entry to be the outward edge of an inward edge at $\ell$-th position in an Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$.
 The condition essentially depends on the number of inward edges and outward edges between the two given positions $i$ and $\ell$.
-Before going into the details, for any two positions i{i} and $\ell$}, with $1\leq\ell<i\leq 2n$, we denote by $\rm{in}_{\ell i}$ (resp., $\rm{out}_{\ell i}$), the number of inward edges (resp., outward edges) $t_{k}$, $\ell<k<i$.
+Before going into the details, for any two positions i{i} and $\ell$}, with $1\leq\ell<i\leq 2n$, we denote by $\rm{in}_{\ell,i}$ (resp., $\rm{out}_{\ell,i}$), the total number of inward (resp., outward) edges among the entries $t_{k}$, $\ell<k<i$, where $t_{k}$ is the $k$-th entry of $E(T)$.
 Consider the tree $T$ given in Fig.
@@
 Similarly, for $i=9,\ell=3$, $\rm{in}_{3,9}=3>2=\rm{out}_{3,9}$ (see Fig.
-4(b)) and for $i=9,\ell=5$, $\rm{in}_{5,9}=1<2=\rm{out}_{5,9}$ (see Fig.
-4(c)).
-Figure 4: An illustration of the number of inward edges (green) (resp., outward edges (blue)) for the edges $t_{i},i=7,9$ and $t_{\ell},\ell=2,3,5$ which are depicted in gray boxes.
+4(b)), for $i=9,\ell=5$, $\rm{in}_{5,9}=1<2=\rm{out}_{5,9}$ (see Fig.
+4(c)), and for $i=7,\ell=4$, $\rm{in}_{4,7}=1=\rm{out}_{4,7}$ (see Fig.
+4(d)).
+Figure 4: (a)-(c)An illustration of the number of inward edges (green) (resp., outward edges (blue)) for the edges $t_{i},i=7,9$ and $t_{\ell},\ell=2,3,5$ which are depicted in gray boxes; (d) An example in which Proposition 1 holds.
 Proposition 1.
@@
   3. (iii)
-    $\rm{in}_{\ell i}=\rm{out}_{\ell i}$, and
+    $\rm{in}_{\ell,i}=\rm{out}_{\ell,i}$, and
   4. (iv)
@@
 That is $t_{2}$ and $t_{4}$ are both candidate inward edges of $t_{7}$.
-By using Proposition 1(iv), $t_{7}$ is the outward edge of $t_{4}$.
+By using Proposition 1(iv), $t_{7}$ is the outward edge of $t_{4}$(see Fig.
+4(d)).
 For $i=9$, we see that $t_{5}$ and $t_{8}$ are the only edges that satisfy Proposition 1(i)-(ii), but $t_{5}$ does not satisfy the Proposition 1(iii).
@@
 In this context $x_{1},\ldots,x_{d}$ represents the inward edges for the substitution operations, while $x_{d+1},\ldots,x_{2d}$ denotes the values to be substituted during these operations.
+The generated Euler strings have distance at most $2d$ as substitution at root and repeated terms in $x$ will be ignored.
 An illustration of such a network is given in Fig.
 6, where $m=5$, $d=3$, $x=1,3,0,5,1,2$, means that the labels of the inward edges (resp., outward edges) of $x_{1}=1,x_{2}=3$ are substituted by $x_{4}=5,x_{5}=1$ (resp., $10,6$).
-The entry $x_{3}=0$ corresponds to the root, and is ignored.
+Since the entry $x_{3}=0$ and $x_{6}=2$ corresponds to the root and its label, resp., and so they are ignored.
 Figure 6: An illustration of a TSd-generative ReLU with the input layer $x=1,3,0,5,1,2$, $E(T)=3,2,7,2,4,9,7,4,9,8$ and output layer $u=E(U)=5,2,7,1,4,9,6,4,9,10$.
@@
 We define the TDd-generative ReLU to be a ReLU neural network that can generate all Euler strings over $\Sigma$ with the tree edit distance at most $2d$ from $E(T)$ by deleting from $E(T)$, the appropriate inward and outward edges of $x=x_{1},x_{2},\ldots,x_{d}$, over $\{0,\ldots,n\}$.
-To output a fixed number of nodes, we assume that $E(T)$ is padded with $2d$ $B$s, $B\gg m$, so that the network can delete the inward and outward edges of $x_{j}\neq 0$ and delete two $B$s corresponding to each $x_{j}=0$.
-We call such a string a padded Euler string.
+The generated Euler strings have distance at most $2d$ as deletion of root and repeated terms in $x$ will be ignored.
+Since the network does not always delete exactly $2d$ elements from the input Euler string, while a neural network must have a fixed number of nodes in each layer, we pad $E(T)$ with $2d$ number of $B$s where $B\gg m$, to ensure a fixed output dimension.
+To delete a vertex, both its inward and outward edges must be deleted.
+Therefore, for each $x_{j}\neq 0$, the network deletes two elements of the input Euler string corresponding to the inward and outward edges of $x_{j}$, and for each $x_{j}=0$, it deletes two padding symbols $B$.
+We call such a string a padded Euler string We call such a string a padded Euler string.
 In this setting, we fix $2n$ output nodes $y=y_{1},\ldots,y_{2n}$ of the network, where $y_{1},\ldots,y_{2n-2d^{\prime}}$, $d^{\prime}\leq d$ is the Euler string of some tree $U$ with the tree edit distance $2d$ from $E(T)$ by deleting $d^{\prime}$ inward and outward edges of non-zero entries of $x$, and the remaining $2d^{\prime}$ entries $y_{2n-2d^{\prime}+1},\ldots,y_{2n}$ are $B$s.
@@
 Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$.
-We define the TId-generative ReLU to be a ReLU neural network with $4d$ input nodes $x=x_{1},x_{2},\ldots,x_{4d}$ over $\{0,\ldots,n\}$, and $2n+2d$ output nodes $u=u_{1},u_{2},\ldots,u_{2n+2d}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance exactly $2d$ from $E(T)$ can be obtained by the insertion of appropriate child nodes $x^{\prime}_{j}$ of $x_{j}$ with inward edges and outward edges of labels $x_{j+3d}$ and $x_{j+3d}+m$, resp., of $x^{\prime}_{j}$, and setting the appropriate ($x_{j+d}$, $x_{j+d}+1$, …, $x_{j+2d}$)-th children of the node $x_{j}$ as the children of $x^{\prime}_{j}$.
+We define the TId-generative ReLU to be a ReLU neural network with $4d$ input nodes $x=x_{1},x_{2},\ldots,x_{4d}$ over $\{0,\ldots,n\}$, and $2n+2d$ output nodes $u=u_{1},u_{2},\ldots,u_{2n+2d}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance exactly $2d$ from $E(T)$ can be obtained by the insertion of appropriate child nodes $x^{\prime}_{j}$ of $x_{j}$ with inward edges and outward edges of labels $x_{j+3d}$ and $x_{j+3d}+m$, resp., of $x^{\prime}_{j}$, and setting the appropriate ($x_{j+d}$, $x_{j+d}+1$, $x_{j+d}+2$, …, $x_{j+2d}$)-th children of the node $x_{j}$ as the children of $x^{\prime}_{j}$.
+It means $x^{\prime}_{j}$ becomes the parent of a sequence of children that starts from $x_{j+d}$ and ends at $x_{j+2d}$.
 In this study, we do not consider inserting new nodes as children of a newly inserted node.
@@
 For convenience, we denote $x_{1},\ldots,x_{d}$ , $x_{1+d},\ldots,x_{2d}$, $x_{1+2d},\ldots,x_{3d}$, and $x_{1+3d},\ldots,x_{4d}$ by $x^{1}$, $x^{2}$, $x^{3}$, and $x^{4}$, respectively.
-Note that some lower and upper bounds of children may not be valid due to the random nature of $x$, and thus need to be refined to perform appropriate insertion operations.
+Note that lower and upper bounds on the number of children must not exceed the total number of children of a node $x_{j}^{1}$.
+Let $D(x_{j}^{1})$ denote the total number of children of a node $x_{j}^{1}$ then we require $x_{j}^{2},x_{j}^{3}\leq D(x_{j}^{1})$.
+Due to the random nature of $x$, some of the lower and upper bounds of children may not be valid, and thus need to be refined to perform appropriate insertion operations.
 Such invalid bounds, their refinements and appropriate insertions are listed in Table 3, where $D(x_{j}^{1})$ denotes the number of children of the node $x_{j}^{1}$.
@@
 Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$.
-We define the TEd-generative ReLU to be a ReLU neural network such that each Euler string over $\Sigma$ with edit distance at most $2d$ from $E(T)$ due to deletion, substitution and insertion operations can be obtained by appropriately choosing an input $x=x_{1},x_{2},\ldots,x_{7d}$ of $7d$ nodes with $x_{j}\in[0,1)$, where $x_{j}$ is of the form $\cdot\Delta, ii$ is an integer and $\Delta$ is a small constant.
+We define the TEd-generative ReLU to be a ReLU neural network such that each Euler string over $\Sigma$ with edit distance at most $2d$ from $E(T)$ due to deletion, substitution and insertion operations can be obtained by appropriately choosing an input $x=x_{1},x_{2},\ldots,x_{7d}$ of $7d$ nodes with $x_{j}\in[0,1)$, where $x_{j}$ is of the form $\cdot\Delta, ii$ is an integer and $\Delta$ is a small constant to set the number of digits to be considered after decimal in $x_{j}$.
+For example, if $\Delta=0.01$, then $x_{j}$ can be any number in [0,1) with two decimal places, i.e., $x_{j}$ cannot take the value 0.011.
+.
 The input $x_{1},x_{2},\ldots,x_{d}$ (resp., $x_{d+1},x_{d+2},\ldots,x_{3d}$ and $x_{3d+1},x_{3d+2},\ldots,x_{7d}$) represents the nodes for deletion (resp., substitution and insertion) operations.
@@
 9 in Appendix 9.
-Figure 9: An illustration of a TEd-generative ReLU for $d=3$, with the input layer x=0.3,0,0.38,0,0.46,0.55,0,0.6,0.88,0.66,0.75,0,0.55,0.87,0.03,0.02x=0.3,0,0.38,0,0.46,0.55,0,0.6,0.88,0.66,0.75,0,0.55,0.87,0.03,0.02, $0.45,0.09,0,0.7,0.5$, padded Euler string $3,2,12,2,4,14,12,4,14,13,B,B,B,B,B,B$ and the output layer $B,B,B,B,5,3,2,6,16,12,4,14,13,15,B,B,B,B$ with $d_{1}=d_{2}=1$.
+Figure 9: An illustration of a TEd-generative ReLU for $d=3$, with the input layer x=x=0.3, 0, 0.38, 0, 0.46, 0.55, 0, 0.6, 0.88, 0.66, 0.75, 0, 0.55, 0.87, 0.03, 0.02, $0.45,0.09,0,0.7,0.5$, the integer conversion of $x$ in 2, 0, 2, 0, 3, 3, 1, 6, 9, 4, 4, 0, 3, 5, 1, 1, 3, 1, 1, 7, 5, padded Euler string $3,2,12,2,4,14,12,4,$ $14,13,B,B,B,B,B,B$ and the output layer $B,B,B,B,5,3,2,6,16,12,4,14,13,15,B,B,B,B$ with $d_{1}=d_{2}=1$.
 By trimming $B$s, we can get the resultant string $E(U)=5,3,2,6,16,12,4,14,13,15$ obtained by deleting, substituting and inserting the indicated nodes.
@@
 We implemented the proposed networks on a machine with an AMD Ryzen 7 4800H, Radeon Graphics processor (2.90 GHz), 16 GB of RAM, and Windows 11 Pro using Python version 3.11.6.
+Note that the proposed $\mathrm{TE}_{d}$-generative ReLU network supports all three edit operations, deletion, substitution, and insertion, whereas the $\mathrm{TI}_{d}$-generative ReLU network supports only insertion operations; however, $\mathrm{TI}_{d}$ has the same order $\mathcal{O}(n^{3})$ as the corresponding $\mathrm{TE}_{d}$ Therefore, for analysis and comparison, we generated trees using the $\mathrm{TI}_{d}$ and $\mathrm{TE}_{d}$ networks for four trees $T_{i}$, $i=1,2,3,4$, shown in Fig.
+10, with $8,10,6,$ and $11$ nodes, respectively, labels from $\Sigma=\{1,2,\ldots,10\}$, and distance bounds $d=2,2,3,$ and $2$.
 Figure 10: Trees used in the computational experiments of TId-generative ReLU and TEd-generative ReLU.
 For each tree $T_{i}$, $i=1,2,3,4$ illustrated in Figs.
-10(a)-(d) with $8,10,6,11$ nodes, labels from $\Sigma=\{1,2,\ldots,10\}$, and $d=2,2,3,2$, resp., we generated all non-isomorphic trees $U$ that have tree edit distance exactly $d$ by using the proposed TId-generative ReLU networks.
-The newly inserted nodes in $T_{1}$ are 7, 7, in $T_{2}$ are 9, 9, in $T_{3}$ are 8, 8, 8, and in $T_{4}$ are 2, 2.
-The computational results such as the number of nodes in each layer of the constructed TId network along with the number of generated trees for each input tree by the network are provided in Table 4.
+10(a)-(d), we supplied the proposed $\mathrm{TI}_{d}$-generative ReLU network with a collection of input sequences $x$ that can cover all possible insertion operations to generate all possible similar trees with distance exactly $d$ due to the insertion operations.
+Duplicate outputs were removed in a post-processing step, yielding the set of distinct trees $U$ whose tree edit distance from $T_{i}$ is exactly $d$.
+In these experiments, two (resp., two; three; and two) new nodes were inserted with labels 7, 7 (resp., 9, 9; 8, 8, 8; and 2, 2) in $T_{1}$ (resp., $T_{2}$; $T_{3}$; and $T_{4}$) to generate similar trees of edit distance exactly $2$ (resp., 2; 3; and 2) from $T_{1}$ (resp., $T_{2}$; $T_{3}$; and $T_{4}$) using the proposed $\mathrm{TI}_{d}$ network.
+The computational results such as the number of nodes in each layer of the constructed TId network along with the number of all distinct trees obtained after removing duplication from the similar trees generated by the proposed network for each input tree are provided in Table 4.
 A summary of these computational results is given below.
-For $T_{i}$, let $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm I}_{i}$ denote the sequence of number of hidden layers, total number of hidden nodes, minimum number of hidden nodes, average number of hidden nodes, and maximum number of hidden nodes, resp., in the TId-generative ReLU network for $T_{i}$.
-From these experiments, we have $(57,32876,6,576.77,11903)^{\rm I}_{1}$, $(57,57324,6,1005.68,24841)^{\rm I}_{2}$, $(57,20556,9,360.63,4500)^{\rm I}_{3}$, $(57,73064,6,1281.82,33860)^{\rm I}_{4}$.
+For $T_{i}$, let $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},\rm{MaxN},RT)^{\rm I}$i denote the sequence of number of hidden layers in TId, total number of hidden nodes in TId, minimum number of hidden nodes in TId, average number of hidden nodes in TId, maximum number of hidden nodes in TId, and running time (sec.) to generate one similar tree using TId-generative ReLU network for $T_{i}$, respectively.
+From these experiments, we have $(57,32876,6,576.77,11903,17.65)^{\rm I}_{1}$, $(57,57324,6,1005.68,24841,46.23)^{\rm I}_{2}$, $(57,20556,9,360.63,$ $4500,7.11)^{\rm I}_{3}$, $(57,73064,6,1281.82,33860,69.84)^{\rm I}_{4}$.
 These values are illustrated in Fig.
@@
 Moreover, the widest layers are significantly larger than the others; for example, in $T_{4}$ with 11 nodes, the widest layer contains 33860 nodes which is over 26.41 times the average layer size (1281.82).
-Additionally, the total number of nodes grows faster than the tree size, e.g., while 11 nodes are roughly twice 6, the corresponding total node count (73064) is 3.56 times that for 6 nodes (20556).
-Figure 11: Information of the architectures of the proposed ReLU-networks for insertion and unified cases and their comparison: (a) Information of $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm I}_{i}$ of the proposed ReLU-networks for insertion case for each tree $T_{i}$, $i=1,2,3,4$; (b) Information of $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm E}_{i}$ of the proposed ReLU-networks for unified case for each tree $T_{i}$, $i=1,2,3,4,5$; (c) A comparison of the information of $(\#\rm{TN},\rm{AvgN},\#\rm{MaxN})_{i}$ of the proposed ReLU-networks for insertion and unified cases for each tree $T_{i}$, $i=1,2,3,4$.
+Additionally, the total number of nodes grows faster than the tree size, e.g., while the number of nodes (11) in $T_{4}$ are roughly 1.8 times the number of nodes (6) in $T_{3}$, the corresponding total node count (73064) in TId for $T_{4}$ is 3.56 times the node count (20556) in TId for $T_{3}$.
+As a result, the running time also increases faster than the tree size.
+Figure 11: Information of the architectures of the proposed ReLU-networks for insertion and unified cases and their comparison: (a) Information of $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},$ $\rm{MaxN},\rm{RT})^{\rm I}$i of the proposed ReLU-networks for insertion case for each tree $T_{i}$, $i=1,2,3,4$; (b) Information of $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},\rm{MaxN},{\rm RT})^{\rm E}$i of the proposed ReLU-networks for unified case for each tree $T_{i}$, $i=1,2,3,4,5$.
+The ranges of L, MinN, AvgN, and RT are much smaller than those of TN and MaxN, which causes their plots to appear compressed and overlap when shown on the same scale; (c) A comparison of the information of $(\rm{TN},\rm{AvgN},\rm{MaxN},\rm{RT})$i of the proposed ReLU-networks for insertion and unified cases for each tree $T_{i}$, $i=1,2,3,4$.
 Similarly, TEd-generative ReLU neural networks were constructed for the trees $T_{i}$, $i=1,2,3,4,5$ given in Figs.
@@
 Table 4: Experimental results for TId-generative ReLU
-Input trees Size of each hidden layer of TId-generative ReLU Number of generated trees
+Input trees Size of each hidden layer of TId-generative ReLU Number of all distinct similar trees obtained by TId
 $T_{1}$, Fig.
@@
 10(d) 48, 148, 1628, 1308, 448, 1766, 808, 408, 29, 1688, 428, 848, 428, 2129, 869, 1436, 33860, 890, 1312, 704, 474, 980, 976, 984, 980, 976, 1000, 984, 976, 996, 984, 974, 972, 984, 974, 974, 972, 1144, 1850, 926, 6, 26, 10, 22, 12, 6, 374, 29, 28, 472, 112, 96, 56, 32, 364, 108, 48 660
-The computational results of these experiments are given in Table 5 which are summarized below.
-For $T_{i}$, let $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm E}_{i}$ denote the sequence of number of hidden layers, total number of hidden nodes, minimum number of hidden nodes, average number of hidden nodes, and maximum number of hidden nodes, resp., in the TEd-generative ReLU network for $T_{i}$.
-For the computational results listed in Table 5 we have $(108,50360,14,466.30,11917)^{\rm E}_{1}$, $(108,82060,14,759.81$, $24859)^{\rm E}_{2}$, $(108,35803,19,331.51,4510)^{\rm E}_{3}$, $(108,101930,14$, $943.80,33880)^{\rm E}_{4}$, $(108,493790,14,4572.13,263350)^{\rm E}_{5}$ for TEd-generative ReLU networks.
+The detailed computational results of these experiments are given in Table 5 which are summarized below.
+For $T_{i}$, let $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},\rm{MaxN},\rm{RT})^{\rm E}$i denote the sequence of number of hidden layers in TEd, total number of hidden nodes, in TEd, minimum number of hidden nodes in TEd, average number of hidden nodes in TEd, maximum number of hidden nodes in TEd, and running time (sec.) to generate one similar tree using TEd-generative ReLU network for $T_{i}$, respectively.
+We have $(108,50360,14,466.30,11917,25.59)^{\rm E}_{1}$, $(108,82060,14,759.81$, $24859,55.23)^{\rm E}_{2}$, $(108,35803,19,$ $331.51,4510,11.36)^{\rm E}_{3}$, $(108,101930,14$, $943.80,33880,$ $85.04)^{\rm E}_{4}$, $(108,493790,14,$ $4572.13,263350,1581.17)^{\rm E}_{5}$ for TEd-generative ReLU networks.
 These values are illustrated in Fig.
@@
 Table 5: Experimental results for TEd-generative ReLU
-Input trees Size of each hidden layer of TEd-generative ReLU Number of generated trees
+Input trees Size of each hidden layer of TEd-generative ReLU Number of all distinct similar trees obtained by TEd
 $T_{1}$, Fig.
@@
 Building on these results, we demonstrated that all similar trees generated through substitution (resp., deletion and insertion) operations can be constructed by ReLU networks of size $\mathcal{O}(dn^{2})$ (resp., $\mathcal{O}(n^{2})$ and $\mathcal{O}(n^{3})$), all with constant depth.
-Finally, we proved that there exists a ReLU network of size $\mathcal{O}(n^{3})$ and constant depth that can generate all trees with a distance at most $d$ to the original tree under combined substitution, deletion, and insertion operations.
+Finally, we proved that there exists a ReLU network of size $\mathcal{O}(n^{3})$ and constant depth capable of generating any tree within distance $d$ from the original tree under combined substitution, deletion, and insertion operations.
 These findings provide a theoretical foundation towards construction of compact generative models and open new directions for efficient tree-structured data generation.